OpenStudy (anonymous):

50 million times 60 trillion = C times 10^M...... Find C and M

jimthompson5910 (jim_thompson5910):

50 million = 50,000,000 = 5 * 10^7 60 trillion = 60,000,000,000,000 = 6 * 10^13 --------------- 50 million times 60 trillion = (5 * 10^7)*(6 * 10^13) = (5*6)*(10^7*10^13) = 30 * 10^(7+13) = 30*10^20 = 3*10^21 So C = 3 and M = 21

OpenStudy (anonymous):

5x10^7*6x10^13=30x10^20=3x10^21